A vegetable fiber is traded in a competitive world market, and the world price is $9 per pound. Unlimited quantities are available for import into the United States at this price. The U.S. domestic supply and demand for various price levels are shown as follows:

a. What is the equation for demand? What is the equation for supply?

b. At a price of $9, what is the price elasticity of demand? What is it at a    price of $12?

c. What is the price elasticity of supply at $9? At $12?

d. In a free market, what will be the U.S. price and level of fiber imports? 

1
Expert's answer
2022-03-22T11:06:16-0400

The equation for demand

i will take two (Q,P) pairs to help me calculate the slope that will be the inverse of the elasticity;

the pairs selected are,

"(34,3), (28,6)"

Find the slope as follows;

"\frac{\Delta P }{\Delta Q} = \frac{ 6-3}{28-34}= -\frac{1}{2}"

To find the demand function we introduce an unknown pair of Q and P

"-\frac{1}{2} = (16,12), (Q,P)"

The equation is therefore;

"P = 28-Q"


The equation for supply

"(6,9), (10,15)"

Find the slope

"\frac{\Delta P }{\Delta Q} = \frac{ 15-6}{10-6}= \frac{3}{2}"

"\frac{3}{2} = (12,18), (Q,P)"

solving this gives the supply equation as;

"P=\frac{3}{2}Q"


c)The price elasticity of demand at $9

"\frac{\Delta Q }{\Delta P}= \frac{22}{9} =2.4"


The price elasticity of demand at $12

"\frac{\Delta Q }{\Delta P}= \frac{16}{12} =1.33"


The price elasticity of supply at $9

"\frac{\Delta Q }{\Delta P}= \frac{6}{9} =0.667"


The price elasticity of supply at $12

"\frac{\Delta Q }{\Delta P}= \frac{8}{12} =0.667"


d) here, we will equate the demand and supply functions obtained above.

 "28-Q =\frac{3}{2}"

"Q= 26.5"

"P = 39.75"